The motion of a mass-spring system with damping is governed by y(t) + 8y (t) + ky(t) = 0; y(0) = 4, and y(0) = 0. Find th
Answer #1

siven that Eruction of motion y[++ 891 € + Kylt) =0, 710) = 4, y10) =0 for k=13 y+8ylt) + 15y =0 AE: mitsm113=0 m=-87 564ч-с, а Ci= 2. 4 - 3 3 - - - 8 3 -(чJa) е (чJ3) е 2 = (2-34) е ек